# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v4

@@ 9.1 Noncommutative Hardy spaces associated to subdiagonal algebras
First, we explain how introduce Hilbert transform associated to subdiagonal algebras of von Neumann algebras. Such algebras were introduced by Arveson in [Arv67] in the case of finite von Neumann algebras and in [Bek15] for the semifinite case. We refer also to [BlL07], [Ji14], [LaX13], [MaW98], [Ran98], [Ran02] and [PiX03] for more information on Hardy spaces associated with subdiagonal algebras of von Neumann algebras. Let $\mathcal{M}$ be a von Neumann algebra endowed with a normal semifinite faithful trace and let $\mathrm{E}\colon\mathcal{M}\to\mathcal{M}$ be a trace preserving conditional expectation onto a von Neumann subalgebra $\mathcal{D}$ of $\mathcal{M}$. Following [Bek15, Definition 2.1 p. 1350], a subdiagonal algebra $\mathrm{H}^{\infty}(\mathcal{M})$ with respect to $\mathrm{E}$ is a weak* closed subalgebra of $\mathcal{M}$ such that
@@
    the trace preserving conditional expectation $\mathrm{E}\colon\mathrm{L}^{\infty}(\mathcal{M})\to\mathrm{L}^{\infty}(\mathcal{M})$ is multiplicative on $\mathrm{H}^{\infty}(\mathcal{M})$, i.e. we have $\mathrm{E}(fg)=\mathrm{E}(f)\mathrm{E}(g)$ for any $f,g\in\mathrm{H}^{\infty}(\mathcal{M})$,
@@
    $\mathrm{H}^{\infty}(\mathcal{M})+(\mathrm{H}^{\infty}(\mathcal{M}))^{*}$ is weak* dense in $\mathrm{L}^{\infty}(\mathcal{M})$.
Here, we warn the reader that $(\mathrm{H}^{\infty}(\mathcal{M}))^{*}$ is not the dual of $\mathrm{H}^{\infty}(\mathcal{M})$, the family of the adjoints of the elements of $\mathrm{H}^{\infty}(\mathcal{M})$. In this case, $\mathcal{D}$ is called the ¡¡diagonal¿¿ of the algebra $\mathrm{H}^{\infty}(\mathcal{M})$. We also define $\mathrm{H}_{0}^{\infty}(\mathcal{M})\overset{\mathrm{def}}{=}\{f\in\mathrm{H}^{\infty}(\mathcal{M}):\mathrm{E}(f)=0\}$. Suppose that $1<p<\infty$. The closures of the subspaces $\mathrm{H}^{\infty}(\mathcal{M})\cap\mathrm{L}^{p}(\mathcal{M})$ and $\mathrm{H}_{0}^{\infty}(\mathcal{M})\cap\mathrm{L}^{p}(\mathcal{M})$ in $\mathrm{L}^{p}(\mathcal{M})$ will be denoted by $\mathrm{H}^{p}(\mathcal{M})$ and $\mathrm{H}^{p}_{0}(\mathcal{M})$. Note that $\mathrm{E}$ extends to a contractive projection $\mathrm{E}\colon\mathrm{L}^{p}(\mathcal{M})\to\mathrm{L}^{p}(\mathcal{M})$ onto the subspace $\mathrm{L}^{p}(\mathcal{D})$ with kernel $\mathrm{H}^{p}_{0}(\mathcal{M})$. By [Bek15, Theorem 4.2 p. 1356], we have a canonical topological direct sum decomposition
@@
(9.5) $F\overset{\mathrm{def}}{=}\mathrm{i}\mathcal{H}+\mathrm{E}\colon\mathrm{L}^{p}(\mathcal{M})\to\mathrm{L}^{p}(\mathcal{M}).$
@@
Suppose that $1<p<\infty$. Assume that the commutator $\mathcal{H},M_{a}]$ is a compact operator acting on $\mathrm{L}^{p}(\mathcal{M})$ for any $a\in\mathcal{A}$ and that the algebra $\mathcal{D}$ is finite-dimensional. Then $(\mathrm{L}^{p}(\mathcal{M}),F)$ is an odd Banach Fredholm module. For any invertible element $a\in\mathcal{A}$, the Toeplitz operator $T_{a}$ is Fredholm and we have
@@
Proof : Since the map $\mathrm{E}\colon\mathrm{L}^{p}(\mathcal{M})\to\mathrm{L}^{p}(\mathcal{M})$ is a projection on the subspace $\mathrm{L}^{p}(\mathcal{D})$, we have with (9.3)
$F^{2}=(\mathrm{i}\mathcal{H}+\mathrm{E})^{2}=\mathrm{Id}.$
@@
$F,M_{a}]=[\mathrm{i}\mathcal{H}+\mathrm{E},M_{a}]=\mathrm{i}[\mathcal{H},M_{a}]+[\mathrm{E},M_{a}]$
@@ 9.3 Fredholm module associated to the discrete Schur-Hilbert transform on $S^{p}_{n}$
Consider the trace preserving conditional expectation $\mathrm{E}\colon\mathrm{M}_{n}\to\mathrm{M}_{n}$ onto the diagonal subalgebra $\mathcal{D}=\ell^{\infty}_{n}$ of $\mathrm{M}_{n}$, where $\mathrm{M}_{n}$ is equipped with its canonical trace. The ¡¡upper triangle subalgebra¿¿ $\mathrm{T}_{n}$ of all upper triangular matrices is a finite subdiagonal algebra of $\mathrm{M}_{n}$. Suppose that $1<p<\infty$. In this context, the Hilbert transform $\mathcal{H}\colon S^{p}_{n}\to S^{p}_{n}$ is the ¡¡discrete Schur-Hilbert transform¿¿ defined as being the Schur multiplier with symbol $\mathrm{i}\,\operatorname{\mathrm{sign}}(i-j)]$. The Riesz projection $P\colon S^{p}_{n}\to S^{p}_{n}$ is the triangular projection. We consider the representation $\pi\colon\mathrm{M}_{n}\to\mathrm{B}(S^{p}_{n})$, $a\mapsto M_{a}$ of the matrix algebra $\mathrm{M}_{n}$, where $M_{a}\colon S^{p}_{n}\to S^{p}_{n}$, $y\mapsto ay$ is the multiplication operator by the matrix $a$. The couple $(S^{p}_{n},F)$ is obviously an odd Banach Fredholm module. The assumptions of Proposition 9.1 are obviously satisfied for the algebra $\mathcal{A}=\mathrm{M}_{n}$. Since $\mathrm{K}_{1}(\mathrm{M}_{n})=0$, we obtain the following result.
@@ 9.4 Banach Fredholm modules on reduced group $\mathrm{C}^{*}$-algebras
Example 9.4
Following [ClR16, p. 14], we say that a group is indicable if it has the group of integers $\mathbb{Z}$ as a quotient, and locally indicable if each of its nontrivial finitely-generated sub-groups is indicable. Locally indicable groups are left-orderable by [ClR16, Corollary 1.51 p. 14].
Very large classes of groups are left-orderable as the following list shows.
Example 9.5
Every free group $\mathbb{F}_{n}$ with an arbitrary number $n$ (possibly infinite) of generators is bi-orderable, as shown in [ClR16, Theorem 3.4 p. 35] or [KaT08, Corollary 7.12 p. 280].
Example 9.6
If an orientable, irreducible7776. A 3-manifold $M$ is irreducible if every smooth 2-sphere in $M$ bounds a 3-dimensional ball in $M$. 3-manifold $M$ has infinite first homology group, then by [ClR16, Theorem 5.3 p. 66] its fundamental group $\pi_{1}(M)$ is locally indicable, hence left-orderable.
Example 9.7
[…]
Example 9.8
Braid groups
[…]
where $n\geqslant 1$ is an integer, are left-orderable, according to [KaT08, Theorem 7.15 p. 283], or [ClR16, Theorem 7.23 p. 103]. Note that $\mathrm{B}_{1}=\{0\}$ and $\mathrm{B}_{2}=\mathbb{Z}$. It is worth noting that the group $\mathrm{B}_{3}$ is isomorphic to the knot group of the trefoil knot by [ClR16, p. 93]. In particular, it is an infinite non-abelian group. For any integer $n\geqslant 3$, the group $\mathrm{B}_{n}$ is not bi-orderable according to [DDRW08, Proposition 1.2 p. 12].
Example 9.9
The infinite braid group
[…]
is left-orderable by [ClR16, Problem 10.11 p. 136].
Example 9.10
According to [ClR16, Theorem 7.10 p. 95], the pure braid group $\mathrm{P}_{n}$ is bi-orderable for any integer $n\geqslant 2$.
Example 9.11
Torsion-free nilpotent groups are bi-orderable, as shown in [DNR16, p. 15]. By the way, we refer to the paper [Sud17] for some information on the K-theory of the reduced group $\mathrm{C}^{*}$-algebra of these groups.
Example 9.12
For each integer $n\geqslant 2$, the Baumslag-Solitar group $\mathrm{BS}(1,n)\overset{\mathrm{def}}{=}\langle a,b\mid a^{-1}ba=b^{n}\rangle$ is bi-orderable by [DNR16, p. 17]. The K-theory groups of its reduced group $\mathrm{C}^{*}$-algebra are described in [PoV18]. Actually, we have $\mathrm{K}_{0}(\mathrm{C}^{*}(\mathrm{BS}(1,n)))=\mathbb{Z}$ and $\mathrm{K}_{1}(\mathrm{C}^{*}(\mathrm{BS}(1,n)))=\mathbb{Z}\oplus\mathbb{Z}/(n-1)\mathbb{Z}$.
Example 9.13
The Thompson group $\mathrm{F}$ is bi-orderable. See Section 9.6 below for more information.
Example 9.14
If $M$ is an orientable closed Seifert fibered manifold with infinite first homology group, then the fundamental group $\pi_{1}(M)$ is left-orderable, according to [ClR16, Theorem 6.12 p. 83].
Example 9.15
The fundamental groups of all closed surfaces, orientable or not, with the exceptions of the projective plane and the Klein bottle are bi-orderable by [RoW01, Theorem 3 p. 314] or [ClR16, Theorem 3.11 p. 36]. If $\mathscr{K}$ is the Klein Bottle then by [ClR16, Problem 1.10 p. 4] its fundamental group $\pi_{1}(\mathscr{K})=\langle x,y\mid xyx^{-1}=y^{-1}\rangle$ is not bi-orderable. However, it is left-orderable by [RoW01, Theorem 3 p. 314]. Finally, the fundamental group $\pi_{1}(\mathrm{P}_{2}(\mathbb{R}))=\mathbb{Z}/2\mathbb{Z}$ of the projective plane $\mathrm{P}_{2}(\mathbb{R})$ is not left-orderable (see also [ClR16, Proposition 6.13 p. 84]). A description of the K-theory groups of the reduced group $\mathrm{C}^{*}$-algebras of these ¡¡surface groups¿¿ is given in [BBV99].
Example 9.16
Let $\Sigma^{b}_{g,n}$ be an orientable connected surface with genus $g$, $b$ boundary components, and $n$ punctures8887. Recall that $P=\{p_{1},\ldots,p_{n}\}$ is a finite subset of distinct points ¡¡removed¿¿ from the surface. $p_{1},\ldots,p_{n}$. If $b>0$ and if the Euler characteristic of $\Sigma^{b}_{g,n}$ satisfies $\chi(\Sigma^{b}_{g,n})<0$ then by [ClR16, Theorem 7.29 p. 107], the mapping class group $\mathrm{Mod}(\Sigma^{b}_{g,n})$ is left-orderable. This group is the group of isotopy classes of orientation-preserving homeomorphisms that restrict to the identity on the boundary of $\Sigma^{b}_{g,n}$.
Example 9.17
Let $M$ be an $n$-dimensional connected piecewise linear manifold (PL manifold). Consider a non-empty closed (n1)(n-1)-dimensional PL submanifold $K$ of $M$. Then by [ClR16, Theorem 8.12 p. 119] the group $\mathrm{PL}_{+}(M,K)$ of orientation-preserving PL homeomorphisms of $M$, fixed on $K$, is locally indicable and therefore left-orderable.
Example 9.18
Let $M$ be a connected smooth $n$-manifold and let $K$ be a non-empty closed (n1)(n-1)-submanifold of $M$. According to [ClR16, Theorem 8.13 p. 119] the group $\mathrm{Diff}^{1}_{+}(M,K)$ of orientation-preserving homeomorphisms of $M$ which are fixed on $K$ and which are continuously differentiable is locally indicable, hence left-orderable.
Example 9.19
The universal cover $\widetilde{\mathrm{PSL}}_{2}(\mathbb{R})$ of the group $\mathrm{PSL}_{2}(\mathbb{R})$ is left-orderable by [ClR16, Example 1.15 p. 5].
Finally, note that both left-orderability and bi-orderability are clearly preserved under taking subgroups and finite products using lexicographic ordering, see [ClR16, Example 1.6 p. 3]
@@
Assume that $G$ is a left-orderable discrete group and endowed with a left-order $\preceq$. In this context, we can introduce the function $\operatorname{\mathrm{sign}}\colon G\to\mathbb{C}$ defined by
@@
Moreover, we can consider the subalgebra $\mathcal{D}\overset{\mathrm{def}}{=}\mathbb{C}1$ of the group von Neumann algebra $\mathrm{VN}(G)$. The linear map $\mathrm{E}\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$, $f\mapsto\tau(f)1$ is a trace preserving normal faithful conditional expectation onto the ¡¡diagonal¿¿ subalgebra $\mathcal{D}$. If $G$ is left-orderable and endowed with a left-order $\preceq$ then it is essentially9998. The proposed definition in $\mathrm{H}^{\infty}(\mathrm{VN}(G))$ is slightly incorrect since the ¡¡noncommutative Fourier series¿¿ of elements in $\mathrm{VN}(G)$ does not converges in general. observed in [GPX22] that
@@
is a subdiagonal algebra whose Hilbert transform $\mathcal{H}\colon\mathrm{L}^{p}(\mathrm{VN}(G))\to\mathrm{L}^{p}(\mathrm{VN}(G))$ coincides with the bounded Fourier multiplier with symbol $\mathrm{i}\operatorname{\mathrm{sign}}$. Actually the boundedness of the last operator was already known for abelian groups [ABG90].
@@
The following result says that the assumptions of Proposition 9.1 are satisfied.
@@
Suppose that $1<p<\infty$. Let $G$ be a left-orderable discrete group endowed with a left-order. The commutator $\mathcal{H},M_{a}]$ is a compact operator acting on the Banach space $\mathrm{L}^{p}(\mathrm{VN}(G))$ for any element $a$ of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\mathrm{red}}(G)$ of the group $G$. Moreover, $(\mathrm{L}^{p}(\mathrm{VN}(G)),F)$ is an odd Banach Fredholm module.
@@
9.5 The index pairing for the reduced group $\mathrm{C}^{*}$-algebra of the braid group $\mathrm{B}_{3}$
[KaT08, Theorem 7.15 p. 283], or [ClR16, Theorem 7.23 p. 103] [DDRW08]
Let $n\geqslant 2$ be an integer. The Dehornoy ordering (also known as the $\sigma$-ordering) of the braid group $\mathrm{B}_{n}$ is a left-ordering introduced in [Deh94] that is defined in terms of representative words of braids as follows. Following [ClR16, p. 95](see also [DDRW08, Definition 1.6 p. 12]), if $1\leqslant i\leqslant n-1$, a word $w$ in the generators $\sigma_{1},\ldots,\sigma_{n-1}$ is called $i$-positive if $w$ contains at least one occurrence of $\sigma_{i}$, no occurrences of $\sigma_{1},\ldots,\sigma_{i-1}$, and every occurrence of $\sigma_{i}$ has positive exponent. Replacing ¡¡positive¿¿ by ¡¡negative¿¿, we obtain the definition of a $i$-negative word.
A braid $\beta\in\mathrm{B}_{n}$ is called $i$-positive (respectively $i$-negative) if it admits a representative word $w$ in the generators $\sigma_{1},\ldots,\sigma_{n-1}$ that is $i$-positive (respectively $i$-negative). The Dehornoy ordering of the braid group $\mathrm{B}_{n}$ is the ordering whose positive elements are the braids $\beta\in\mathrm{B}_{n}$ that are $i$-positive for some $i$.
Recall that the non-abelian group $\mathrm{B}_{3}$ is defined by

### v5

@@ 9.1 Noncommutative Hardy spaces associated to subdiagonal algebras
First, we explain how introduce Hilbert transform associated to subdiagonal algebras of von Neumann algebras. Such algebras were introduced by Arveson in [Arv67] in the case of finite von Neumann algebras and in [Bek15] for the semifinite case. We refer also to [BlL07], [Ji14], [LaX13], [MaW98], [Ran98], [Ran02] and [PiX03] for more information on Hardy spaces associated with subdiagonal algebras of von Neumann algebras. Let $\mathcal{M}$ be a von Neumann algebra endowed with a normal semifinite faithful trace and let $\mathbb{E}\colon\mathcal{M}\to\mathcal{M}$ be a trace preserving conditional expectation onto a von Neumann subalgebra $\mathcal{D}$ of $\mathcal{M}$. Following [Bek15, Definition 2.1 p. 1350], a subdiagonal algebra $\mathrm{H}^{\infty}(\mathcal{M})$ with respect to $\mathbb{E}$ is a weak* closed subalgebra of $\mathcal{M}$ such that
@@
    the trace preserving conditional expectation $\mathbb{E}\colon\mathrm{L}^{\infty}(\mathcal{M})\to\mathrm{L}^{\infty}(\mathcal{M})$ is multiplicative on $\mathrm{H}^{\infty}(\mathcal{M})$, i.e. we have $\mathbb{E}(fg)=\mathbb{E}(f)\mathbb{E}(g)$ for any $f,g\in\mathrm{H}^{\infty}(\mathcal{M})$,
@@
    the sum $\mathrm{H}^{\infty}(\mathcal{M})+(\mathrm{H}^{\infty}(\mathcal{M}))^{*}$ is weak* dense in $\mathrm{L}^{\infty}(\mathcal{M})$.
Here, we warn the reader that $(\mathrm{H}^{\infty}(\mathcal{M}))^{*}$ is not the dual of $\mathrm{H}^{\infty}(\mathcal{M})$, the family of the adjoints of the elements of $\mathrm{H}^{\infty}(\mathcal{M})$. In this case, $\mathcal{D}$ is called the ¡¡diagonal¿¿ of the algebra $\mathrm{H}^{\infty}(\mathcal{M})$. We also define $\mathrm{H}_{0}^{\infty}(\mathcal{M})\overset{\mathrm{def}}{=}\{f\in\mathrm{H}^{\infty}(\mathcal{M}):\mathbb{E}(f)=0\}$.
Suppose that $1<p<\infty$. The closures of the subspaces $\mathrm{H}^{\infty}(\mathcal{M})\cap\mathrm{L}^{p}(\mathcal{M})$ and $\mathrm{H}_{0}^{\infty}(\mathcal{M})\cap\mathrm{L}^{p}(\mathcal{M})$ in $\mathrm{L}^{p}(\mathcal{M})$ will be denoted by $\mathrm{H}^{p}(\mathcal{M})$ and $\mathrm{H}^{p}_{0}(\mathcal{M})$. Note that $\mathbb{E}$ extends to a contractive projection $\mathbb{E}\colon\mathrm{L}^{p}(\mathcal{M})\to\mathrm{L}^{p}(\mathcal{M})$ onto the subspace $\mathrm{L}^{p}(\mathcal{D})$ with kernel $\mathrm{H}^{p}_{0}(\mathcal{M})$. By [Bek15, Theorem 4.2 p. 1356], we have a canonical topological direct sum decomposition
@@
(9.5) $F\overset{\mathrm{def}}{=}\mathrm{i}\mathcal{H}+\mathbb{E}\colon\mathrm{L}^{p}(\mathcal{M})\to\mathrm{L}^{p}(\mathcal{M}).$
@@
Suppose that $1<p<\infty$. Assume that the commutator $\mathcal{H},M_{a}]$ is a compact operator acting on $\mathrm{L}^{p}(\mathcal{M})$ for any $a\in\mathcal{A}$ and that the algebra $\mathcal{D}$ is finite-dimensional. Then $(\mathrm{L}^{p}(\mathcal{M}),\pi,F)$ is an odd Banach Fredholm module. For any invertible element $a\in\mathcal{A}$, the Toeplitz operator $T_{a}$ is Fredholm and we have
@@
Proof : Since the map $\mathbb{E}\colon\mathrm{L}^{p}(\mathcal{M})\to\mathrm{L}^{p}(\mathcal{M})$ is a projection on the subspace $\mathrm{L}^{p}(\mathcal{D})$, we have with (9.3)
$F^{2}=(\mathrm{i}\mathcal{H}+\mathbb{E})^{2}=\mathrm{Id}.$
@@
$F,M_{a}]=[\mathrm{i}\mathcal{H}+\mathbb{E},M_{a}]=\mathrm{i}[\mathcal{H},M_{a}]+[\mathbb{E},M_{a}]$
@@ 9.3 Fredholm module associated to the discrete Schur-Hilbert transform on $S^{p}_{n}$
Consider the trace preserving conditional expectation $\mathrm{E}\colon\mathrm{M}_{n}\to\mathrm{M}_{n}$ onto the diagonal subalgebra $\mathcal{D}=\ell^{\infty}_{n}$ of $\mathrm{M}_{n}$, where $\mathrm{M}_{n}$ is equipped with its canonical trace. The ¡¡upper triangle subalgebra¿¿ $\mathrm{T}_{n}$ of all upper triangular matrices is a finite subdiagonal algebra of $\mathrm{M}_{n}$. Suppose that $1<p<\infty$. In this context, the Hilbert transform $\mathcal{H}\colon S^{p}_{n}\to S^{p}_{n}$ is the ¡¡discrete Schur-Hilbert transform¿¿ defined as being the Schur multiplier with symbol $\mathrm{i}\,\mathrm{sign}(i-j)]$. The Riesz projection $P\colon S^{p}_{n}\to S^{p}_{n}$ is the triangular projection. We consider the representation $\pi\colon\mathrm{M}_{n}\to\mathrm{B}(S^{p}_{n})$, $a\mapsto M_{a}$ of the matrix algebra $\mathrm{M}_{n}$, where $M_{a}\colon S^{p}_{n}\to S^{p}_{n}$, $y\mapsto ay$ is the multiplication operator by the matrix $a$. The couple $(S^{p}_{n},F)$ is obviously an odd Banach Fredholm module. The assumptions of Proposition 9.1 are obviously satisfied for the algebra $\mathcal{A}=\mathrm{M}_{n}$. Since $\mathrm{K}_{1}(\mathrm{M}_{n})=0$, we obtain the following result.
@@ 9.4 Banach Fredholm modules on reduced group $\mathrm{C}^{*}$-algebras
Assume that $G$ is a left-orderable discrete group and endowed with a left-order $\preceq$. In this context, we can introduce the function $\mathrm{sign}\colon G\to\mathbb{C}$ defined by
@@
Moreover, we can consider the subalgebra $\mathcal{D}\overset{\mathrm{def}}{=}\mathbb{C}1$ of the group von Neumann algebra $\mathrm{VN}(G)$. The linear map $\mathbb{E}\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$, $f\mapsto\tau(f)1$ is a trace preserving normal faithful conditional expectation onto the ¡¡diagonal¿¿ subalgebra $\mathcal{D}$. If $G$ is left-orderable and endowed with a left-order $\preceq$ then it is essentially7776. The proposed definition in $\mathrm{H}^{\infty}(\mathrm{VN}(G))$ is slightly incorrect since the ¡¡noncommutative Fourier series¿¿ of elements in $\mathrm{VN}(G)$ does not converges in general. observed in [GPX22] that
@@
is a subdiagonal algebra whose Hilbert transform $\mathcal{H}\colon\mathrm{L}^{p}(\mathrm{VN}(G))\to\mathrm{L}^{p}(\mathrm{VN}(G))$ coincides with the bounded Fourier multiplier with symbol $\mathrm{i}\,\mathrm{sign}$. Actually the boundedness of the last operator was already known for abelian groups [ABG90].
@@
The following result says that the assumptions of Proposition 9.1 are satisfied under additional condition.
@@
Suppose that $1<p<\infty$. Let $G$ be a left-orderable discrete group endowed with a left-order such that the set $\{t\in G:\operatorname{\mathrm{sgn}}(st)\not=\operatorname{\mathrm{sgn}}(t)\}$ is finite for any $s\in G$. The commutator $\mathcal{H},M_{a}]$ is a compact operator acting on the Banach space $\mathrm{L}^{p}(\mathrm{VN}(G))$ for any element $a$ of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\mathrm{red}}(G)$ of the group $G$. Moreover, $(\mathrm{L}^{p}(\mathrm{VN}(G)),\pi,F)$ is an odd Banach Fredholm module.
@@
9.5 Banach Fredholm module associated to the free Hilbert transform
Let $\mathbb{F}_{\infty}$ be the free group with a countable sequence $g_{1},g_{2},g_{3}\ldots$ of generators. We only use reduced words. For any integer $n\geqslant 1$, consider the orthogonal projections $L_{n}^{+}\colon\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))\to\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))$ and $L_{n}^{-}\colon\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))\to\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))$ defined by